ganeshie8 (ganeshie8):

[SOLVED by @adxpoi ] show that there are \(\dfrac{n(n+1)(2n+1)}{6}\) triangles made from the vertices of a regular \(2n+1\)-gon such that the center of inscribed circle lies interior to the triangle

Parth (parthkohli):

So does that mean we want acute triangles?

ganeshie8 (ganeshie8):

Precicely! https://i.gyazo.com/201ec61d12a823629c7fee212b70b2ee.png

OpenStudy (empty):

Each n adds a new square that much I know. :P

ganeshie8 (ganeshie8):

Haha, suppose we don't know the formula..

OpenStudy (empty):

So making them odd sided removes any conflict with a point lying on the boundary of our triangle that's helpful at least.

ganeshie8 (ganeshie8):

I think this way of counting also gives a new way to derive formula for sum of first n squares

Parth (parthkohli):

So first of all, each interior angle in a regular 2n+1-gon is\[\frac{180(2n-1 )}{2n+1} \]which is greater than \(90 \) when \(n\ge \frac{3}{2}\) so basically it's always obtuse except for triangles. So one thing we're sure about is that no interior angle can be a part of the triangles we make. That means that the vertices cannot be adjacent to each other.

Parth (parthkohli):

Now I think it boils down to counting and stuff.

ganeshie8 (ganeshie8):

right, another trivial observation : If you fix the first vertex of triangle, then the other two vertices must chosen diametrically opposite such that the triangle captures the center

ganeshie8 (ganeshie8):